OpenStudy (anonymous):

Write an equation in slope-intercept form of the line through points S(–10, –3) and T(–1, 1). (1 point) y=-4/9x + 13/9 y=4/9x - 13/9 y=-4/9x - 13/9 y=4/9x + 13/9

OpenStudy (amistre64):

subtract the points and stack y/x for the slope

OpenStudy (amistre64):

-4/-9 is what I get for the slope; so its either the 2nd or 4th one

OpenStudy (amistre64):

y = mx -mPx + Py is a quick way to develop the slope intercept form; just pick a point they gave in plug in to solve

OpenStudy (anonymous):

hmmm

OpenStudy (amistre64):

y = mx -mPx + Py y = (4/9)x -4(-1)/9 + 1

OpenStudy (amistre64):

looks to be the last one

OpenStudy (amistre64):

the quickest way tho; is just plug the points into the equations and see which one makes it true since they already give you a correct answer in there somewhere

OpenStudy (anonymous):

that would be easier lol thankyou so much :) i have alot of trouble with these <3